# Large Language Models and Code Security: A Systematic Literature Review

### v1

@@ V-C Comparison with Other Approaches
LLMs have shown great promise in vulnerability detection, but it is essential to compare their performance against traditional approaches, such as SATs and other ML models, used in this field for many years. Some studies suggest that LLMs outperform state-of-the-art SATs and ML models in detecting vulnerabilities, while others disagree, highlighting a divide in current research on their comparative effectiveness. Table IV summarizes the reference papers and the approaches used in each study this comparison, including both SAT-based and ML-based methods.
TABLE IV: SAT-based and ML-based approaches used as baselines in the studies.
References Approach
@@
Yu et al. [58] Bandit, CodeQL, CppCheck, Semgrep, SonarQube
Zhang et al. [68] Bugram, CFGNN
Bakhshandeh et al. [69] Bandit, Semgrep, SonarQube
Mohajer et al. [44] Infer
Purba et al. [60] Checkmarx, Flawfinder, RATS
Thapa et al. [45] BiLSTM, BiGRU
Zhou et al. [70] Bandit, CodeChecker, CodeQL, Contrast Codesec Scan, Cppcheck, DevSkim, Dlint, Flawfinder, Graudit, Horusec, Infer, Insider, Semgrep, SonarQube, SpotBugs
Tamberg et al. [71] CodeQL, SpotBugs
Zhang et al. [46] VulTeller, Transformer
Noever et al. [72] Snyk, Fortify
Wen et al. [47] Devign, Reveal
Omar et al. [48] SySeVR
Chen et al. [65] Conkas, Oyente, Smartcheck, Honeybadger, Maian, Securify, Slither, Osiris, Smartian, Confuzzius, Sailfish, Mythril, Solhint, AChecker
Gao et al. [67] LineVul, CodeXGLUE, Cppcheck, Flawfinder, BinAbsInspector
@@
Wen et al. [47] Cppcheck, Flawfinder, RATS, Semgrep, Devign, Reveal
Hu et al. [53] Slither
Li et al. [54] CodeQL
Ni et al. [73] Devign, ReVeal, IVDetect, LineVul, SVulD
Li et al. [55] UBITect
@@
Ozturk et al. [74] Betterscan CE, OWASP WAP, phpcs-security-audit (PHPcs), Pixy, Progpilot, Psalm, RATS, RIPS, SonarQube, Visual Code Grepper (VCG), Horusec
@@
According to Yu et al. [58], traditional code analysis tools, despite their long-standing use, often struggle with issues such as high false positive rates, limited input ranges, and poor scalability. In contrast, LLMs, particularly GPT-based models provide higher detection rates, compared to contemporary SATs like CppCheck [75], Bandit [76], CodeQL [77], SonarQube [78], and Semgrep [79].
For example, in the experiment conducted by Mohajer et al. [44], ChatGPT outperforms Infer [80], offering greater effectiveness in vulnerability detection and code analysis. ChatGPT surpassed Infer, improving precision by around 13% for Null Dereference bugs and 43% for Resource Leak bugs. Similarly, in [72] GPT-4 outperformed Fortify [81] and Snyk [82] detecting roughly four times more vulnerabilities. In another study [54], researchers combined SAT CodeQL with LLMs, achieving a 35% increase in vulnerability detection compared to using CodeQL alone. Some SATs, such as Cppcheck, Flawfinder, RATS, and Semgrep, seem to perform poorly with respect to LLMs. As observed by Wen et al. [51], the reason seems to lie in the fact that these tools have a clear focus on specific vulnerabilities.
While some researchers find SATs to perform poorly in vulnerability detection, other studies suggest that they outperform LLMs in certain aspects, such as reducing false positives. For instance, Purba et al. [60] found that LLMs do not perform as well as SATs, such as Flawfinder [83], RATS [84], and Checkmarx [85] in detecting software vulnerabilities, primarily due to high false positive rates. They therefore suggested using LLMs in combination with SATs for improved performance.
@@
Zhou et al. [70] found that LLMs can detect up to 100% of code vulnerabilities, but they often produce false positive results. In contrast, although SATs achieve lower vulnerability detection rates of up to 44.4%, they also exhibit lower false positive rates. The authors concluded that their extensive comparison of 15 different SATs, including CodeQL, Bandit, and Infer, against 12 state-of-the-art LLMs, revealed significant differences in detecting software vulnerabilities across repositories of three programming languages: Java, C, and Python. Similarly, Tamberg et al. [71] highlighted that LLMs can identify more vulnerabilities than traditional SATs like CodeQL and SpotBugs, but they are prone to non-deterministic behavior and higher false positive rates. Despite these challenges, LLMs adaptability and contextual understanding make them effective tools for vulnerability detection, especially in identifying a wider range of security issues.
@@
Bakhshandeh et al. [69] investigated the applicability of ChatGPT (in the form of GPT-3.5), for vulnerability detection on Python code. They compared the performance of ChatGPT against three SATs: Bandit, Semgrep, SonarQube. In the comparison, ChatGPT was used not only as a stand-alone tool but also as an assistant module for the SATs. The study indicated that while LLMs do not outperform traditional SATs, they show potential when combined with SATs to reduce false positives and negatives.
Further research by Chen et al. [65] explored the performance of ChatGPT against 14 smart contract vulnerability detection tools, including Conkas [86], Oyente [87], and Slither [88]. Their results showed that ChatGPT performed worse than these tools on 71.4% (5/7) of vulnerabilities, showing high efficiency but more detection failures than 71.4% (10/14) of the tools tested. However, in a related experiment by Hu et al. [53], the LLM framework outperformed the smart contract detection tool Slither.
Other researchers turned to ML-based approaches as baselines to assess the effectiveness of LLMs in detecting code vulnerabilities. For instance, Gao et al. [67] employed three Deep Learning (DL) models (i.e., LineVul [89], CodeXGLUE [90], and VulBERTa [91]), along with three rule-based SATs (i.e., Cppcheck, Flawfinder, and BinAbsInspector [92]) as baselines in their research. Their results suggest that GPT-3.5 and GPT-4 outperform SATs and other DL models, such as VulBERTa and CodeXGLUE in vulnerability detection. However, open-access LLM models, (e.g., Llama 2 and Vicuna), do not achieve the same performance, even when trained on the same datasets. Overall, there seems to be a significant gap in performance, between advanced models such as GPT-3.5 and large-scale open-access alternatives. Similarly, ChatGPT has demonstrated significant advantages over ML-based approaches [68], such as Bugram [93] and CFGNN [94], in vulnerability detection tasks.
@@
Furthermore, the framework built upon ChatGPT by Wen et al. [47], was able to capture more vulnerabilities than the tools Devign [97] and Reveal [98]. Similarly, SySeVR, an ML-based model, failed to match the performance of LLMs in vulnerability detection, confirming LLMs superiority [48].
@@
In another study, Ni et al. [73] compared sequence-based and graph-based models for vulnerability detection. They found that sequence-based models, such as LineVul and SVulD, generally outperformed graph-based models like Devign, ReVeal, and IVDetect in accuracy and stability. However, graph-based models showed strengths in certain types of vulnerabilities like API Abuse and Security Features. The study also noted that LLMs are not as effective for direct use in vulnerability detection, and that their performance is highly dependent on prompt design.
@@ V-D Factors Influencing LLMs Performance
One of the issues arises when resolved vulnerabilities remain in the code, often in the form of comments and annotations left by the developers. LLMs may mistakenly interpret these resolved issues as existing vulnerabilities, which can lower their ability to accurately detect existing security vulnerabilities. This implies that there is a need for developers to standardize code commenting practices, to avoid confusion and improve LLMs performance in detecting security vulnerabilities. Additionally, when handling longer pieces of code, LLMs often struggle to identify smaller, more subtle security defects [58].
@@
Some authors have focused their research on identifying and addressing the factors that influence LLMs performance in code analysis tasks, aiming to enhance their accuracy and reliability. Specifically, Wang et al. [101] investigated the impact of naming on LLMs code analysis. In their experiment, they created datasets containing code with nonsensical or misleading variables, methods, and function names. Using CodeBERT for the analysis, they demonstrated that naming significantly affects LLMs ability to detect vulnerabilities, indicating that LLMs rely heavily on clear and meaningful names.
Fang et al. [102] identified a gap in research, concerning how effective LLMs are at analyzing code, particularly when dealing with confusing, obfuscated, or poorly readable code. The authors first conducted experiments on non-obfuscated code, and found that GPT-3.5 and GPT-4 were more likely than smaller models to generate accurate and detailed explanations for input code. On the other hand, obfuscated code impacted the ability of LLMs to generate explanations. GPT-3.5 and GPT-4 both had a drop in analysis accuracy, while smaller models were unable to handle obfuscated code. In conclusion, LLMs from the GPT series performed impressively in code analysis tasks, but they encountered limitations when it comes to obfuscated code, indicating that clarity of the code significantly impacts LLMs analysis performance.
@@ V-E LLMs Fine-tuning for Vulnerability Detection
To enhance the performance of LLMs on vulnerability detection tasks, several researchers have explored fine-tuning techniques. This process, illustrated in Figure 3, involves adapting pre-trained models to security-related tasks, by further training them on security-specific dataset.
@@
After evaluating LLMs for vulnerability detection, the authors of [61] fine-tuned one of the models that performed poorly on vulnerability detection tasks, Meta-Llama-3-8B-Intruct, to verify the impact of fine-tuning on the domain-specific performance. The results showed that fine-tuning can have a positive impact on vulnerability detection tasks, but challenges in identifying exact vulnerabilities remain.
Fine-tuning was also explored by Ding et al. [62]. Despite applying advanced techniques, such as class weights and contrastive learning, LLMs struggled with detecting realistic vulnerabilities.
@@ VI-A Uses of LLMs for Fixing Vulnerabilities
Berabi et al. [113] proposed a new method for fixing code vulnerabilities using LLMs, which significantly enhances their ability to fix vulnerabilities. To increase the model performance, their approach combines static code analysis with code reduction, a technique that involves selecting only the code that contains a defect and the relevant context for interpreting it, and feeding both to the LLM. Results showed that this approach improved the efficiency of the model, by allowing it to understand and focus only on the sections that require fixing.
@@ VI-B LLMs Fine-tuning for Fixing Vulnerabilities
Wu et al. [117] evaluated the performance of different LLMs in fixing Java vulnerabilities, with and without fine-tuning, on two benchmarks they proposed. Across 5 non-fine-tuned LLMs, the best-performing model, Codex, fixed 20.4% of vulnerabilities. Concerning the fine-tuning experiment, the authors could fine-tune 4 models out of 5, since OpenAI does not provide a public API for fine-tuning Codex. Among the fine-tuned models, the best performer (i.e., fine-tuned InCoder), fixed only 18% of vulnerabilities, a small improvement of 8% over the non-fine-tuned version. Besides, these 4 fine-tuned models could only fix vulnerabilities that required simple changes, such as deleting statements, renaming variables, or renaming methods. The models struggled with more complex vulnerabilities, such as CWE-325: Missing Cryptographic Step, and CWE-444: HTTP Request/Response Smuggling.
@@ VII LLMs for Detecting and Fixing Vulnerabilities in a Unified Process
In Section V and Section VI, we have discussed the use of LLMs as tools for handling separately vulnerability detection and vulnerability fixing. In this section, we explore an integrated approach where LLMs are utilized to detect and fix security vulnerabilities in a unified process. Additionally, it examines the capability of LLMs to identify vulnerabilities in the code they generate, and automatically resolve them. Last, the section covers the use of fine-tuning for enhancing LLMs performance, in a unified pipeline for vulnerability detection and fixing.

### v2

@@ V-C Comparison of LLMs with Traditional Vulnerability Detection Approaches
LLMs have shown great promise in vulnerability detection, but it is essential to compare their performance against traditional approaches, such as SATs and other ML models, used in this field for many years. Some studies suggest that LLMs outperform state-of-the-art SATs and ML models in detecting vulnerabilities, while others disagree, highlighting a divide in current research on their comparative effectiveness.
Table V, and Table VI provide an overview of the different LLM models and traditional analysis tools that have been used in each study, for comparing performance. Table V lists the papers where LLMs showed better vulnerability detection capabilities than traditional approaches. In contrast, Table VI highlights the papers where traditional approaches performed better, identifying a higher percentage of vulnerabilities. From the sheer number of papers included in each table, it is clear that, in most literature works, LLMs showed greater potentiality in discovering security vulnerabilities.
TABLE V: Studies showing that LLMs outperformed traditional approaches, based on the number of vulnerabilities detected.
References LLMs Traditional Approaches
@@
Yu et al. [54] GPT (3.5, 4, 4 Turbo), Llama 2 (7B, 70B), Gemini Pro Bandit, CodeQL, CppCheck, Semgrep, SonarQube
Zhang et al. [68] ChatGPT-4 Bugram, CFGNN
Mohajer et al. [44] ChatGPT (3.5 Turbo, 4) Infer
Thapa et al. [45] GPT (2, J, MegatronGPT-2), BERT, MegatronBERT, DistilBERT, RoBERTa, CodeBERT BiLSTM, BiGRU
Noever et al. [69] GPT-4 Snyk, Fortify
Wen et al. [47] Code-llama-7b, GPT-3.5 Devign, Reveal
Omar et al. [48] GPT-2 SySeVR
Gao et al. [63] GPT (3.5, 4), Llama 2 (7b-chat-hf, 13b-chat-hf, 70b-chat-hf), Vicuna (7b-v1.5, 7b-v1.5-16k, 13b-v1.5, 13b-v1.5-16k, 33b-v1.3), ChatGLM2-6b, Baichuan2-13B-Chat, Internlm-20b-chat, CodeLlama-34b-Instruct, Falcon-40b-instruct, Platypus2-70B-instruct LineVul, CodeXGLUE, Cppcheck, Flawfinder, BinAbsInspector
@@
Wen et al. [51] CodeBERT, CodeT5, UniXcoder, EPVD, LineVul, SVulD Cppcheck, Flawfinder, RATS, Semgrep, Devign, Reveal
Hu et al. [65] GPT-4 Slither
@@
Ozturk et al. [70] ChatGPT-3.5 Betterscan CE, OWASP WAP, phpcs-security-audit (PHPcs), Pixy, Progpilot, Psalm, RATS, RIPS, SonarQube, Visual Code Grepper (VCG), Horusec
Purba et al. [56]* GPT-3.5-Turbo, Davinci, CodeGen-2B-multi model Checkmarx, Flawfinder, RATS
Zhou et al. [66]* CodeBERT, G-BERT, CodeT5, UniXcoder, StarCoder, CodeLlama, Mistral, DSCoder, StarCoder2, CodeQwen, LLama3, Phi3 Bandit, CodeChecker, CodeQL, Contrast Codesec Scan, Cppcheck, DevSkim, Dlint, Flawfinder, Graudit, Horusec, Infer, Insider, Semgrep, SonarQube, SpotBugs
Tamberg et al. [67]* GPT (4, 4-turbo), Claude 3 Opus CodeQL, SpotBugs
Note: Studies marked with * indicate cases where LLMs detected more vulnerabilities than traditional approaches but showed significantly higher false positive rates.
TABLE VI: Studies showing that traditional approaches outperformed LLMs, based on the number of vulnerabilities detected.
References LLMs Traditional Approaches
Bakhshandeh et al. [71] GPT-3.5-Turbo Bandit, Semgrep, SonarQube
Zhang et al. [46] GPT (3.5, 4), Llama 2, CodeLlama, WizardCoder, CodeBERT, GraphCodeBERT, PLBART, CodeT5, CodeGen VulTeller, Transformer
Chen et al. [61] GPT (3.5, 4) Conkas, Oyente, Smartcheck, Honeybadger, Maian, Securify, Slither, Osiris, Smartian, Confuzzius, Sailfish, Mythril, Solhint, AChecker
Ni et al. [72] ChatGPT-3.5 Turbo Devign, ReVeal, IVDetect, LineVul, SVulD
@@
According to Yu et al. [54], traditional code analysis tools, despite their long-standing use, often struggle with issues such as high false positive rates, limited input ranges, and poor scalability. In contrast, LLMs, particularly GPT-based models, provide higher detection rates, compared to contemporary SATs. While most studies suggest that LLMs outperform traditional approaches in vulnerability detection, LLMs often exhibit also high false positive rates. For instance, Purba et al.[56] found that LLMs achieved higher detection rates than SATs, but exhibited false positive rates exceeding 60%, compared to rates under 45% for SATs like Flawfinder [73] and Checkmarx [74].
@@
Building on this, Zhou et al. [66] found that LLMs can detect up to 100% of code vulnerabilities, but they often produce false positive results. In contrast, although SATs achieve lower vulnerability detection rates of up to 44.4%, they also exhibit lower false positive rates. The authors concluded that their extensive comparison of 15 different SATs, against 12 state-of-the-art LLMs, revealed significant differences in detecting software vulnerabilities across repositories of three programming languages: Java, C, and Python. Similarly, Tamberg et al. [67] highlighted that LLMs can identify more vulnerabilities than traditional SATs, but they are prone to non-deterministic behavior and higher false positive rates. Despite these challenges, LLMs adaptability and contextual understanding make them effective tools for vulnerability detection, especially in identifying a wider range of security issues.
@@
Bakhshandeh et al. [71] investigated the applicability of ChatGPT (in the form of GPT-3.5), for vulnerability detection on Python code. They compared the performance of ChatGPT against three SATs In the comparison, ChatGPT was used not only as a stand-alone tool but also as an assistant module for the SATs. The study indicated that while LLMs do not outperform traditional SATs, they show potential when combined with SATs to reduce false positives and negatives.
Further research by Chen et al. [61] explored the performance of ChatGPT against 14 smart contract vulnerability detection tools. Their results showed that ChatGPT performed worse than these tools on 71.4% (5/7) of vulnerabilities, showing high efficiency but more detection failures than 71.4% (10/14) of the tools tested. However, in a related experiment by Hu et al. [65], the LLM framework outperformed the smart contract detection tool Slither [75].
Other researchers turned to ML-based approaches as baselines to assess the effectiveness of LLMs in detecting code vulnerabilities. For instance, Gao et al. [63] employed three Deep Learning (DL) models, along with three rule-based SATs as baselines in their research. Their results suggest that GPT models outperform SATs and other DL models, such as VulBERTa and CodeXGLUE in vulnerability detection. However, open-access LLM models, (e.g., Llama 2 and Vicuna), do not achieve the same performance, even when trained on the same datasets. Overall, there seems to be a significant gap in performance, between advanced models such as GPT-3.5 and large-scale open-access alternatives. Similarly, Zhang et al. [68] concluded that ChatGPT has demonstrated significant advantages over ML-based approaches, such as Bugram [76] and CFGNN [77], in vulnerability detection tasks.
@@
Furthermore, the framework built upon ChatGPT by Wen et al. [47], was able to capture more vulnerabilities than the ML tools Devign [80] and Reveal [81]. Similarly, SySeVR, an ML-based model, failed to match the performance of LLMs in vulnerability detection, confirming LLMs superiority [48].
@@
In another study, Ni et al. [72] compared sequence-based and graph-based models for vulnerability detection. They found that sequence-based models, such as LineVul [84] and SVulD [85], generally outperformed graph-based models like Devign, ReVeal, and IVDetect [86] in accuracy and stability. However, graph-based models showed strengths in certain types of vulnerabilities like API Abuse and Security Features. The study also noted that LLMs are not as effective for direct use in vulnerability detection, and that their performance is highly dependent on prompt design.
Some studies indicate that combining LLMs with SATs can improve vulnerability detection rates. For instance, Li et al. [87] observed that LLMs often struggle with complex code reasoning for vulnerability detection and proposed IRIS, a method that combines LLMs with SATs. IRIS uses LLMs to find specific taint specifications related to a vulnerability class (CWE) in third-party library APIs within a project. IRIS improves these specifications with CodeQL to find security vulnerabilities. They evaluated IRIS on a dataset they created with 120 security vulnerabilities found in real-world Java projects, and validated manually. Results showed that IRIS achieved a 35% increase in vulnerability detection compared to using CodeQL alone, while also reducing the number of false positives by 80%. In summary, the combination of LLMs and SATs has enhanced the analysis of vulnerabilities.
Similarly, Li et al. [88] proposed LLift, a framework that combines LLMs with static analysis (using UBITect, a UBI discovery tool), to improve bug detection in large and complex codebases, with a focus on Use Before Initialization (UBI) bugs in the Linux kernel. First, the pipeline uses UBITect to perform a path-insensitive static analysis for scalability, followed by symbolic execution for precision. Then, in the second phase of the pipeline, an LLM is used to address the undecided cases that UBITect cannot handle. The results showed that LLift reduces the complexity of path-sensitive analysis, improving bug detection accuracy without a high rate of false positives. Additionally, LLift was able to identify 4 UBI bugs in the Linux kernel that UBITect could not confirm.
@@ V-D Factors Influencing LLMs Performance
One of the issues arises when resolved vulnerabilities remain in the code, often in the form of comments and annotations left by the developers. LLMs may mistakenly interpret these resolved issues as existing vulnerabilities, which can lower their ability to detect existing security vulnerabilities. This implies that there is a need for developers to standardize code commenting practices, to avoid confusion and improve LLMs performance in detecting security vulnerabilities. Additionally, when handling longer pieces of code, LLMs often struggle to identify smaller, more subtle security defects [54].
@@
Some authors have focused their research on identifying and addressing the factors that influence LLMs performance in code analysis tasks, aiming to enhance their accuracy and reliability. Specifically, Wang et al. [89] investigated the impact of naming on LLMs code analysis. In their experiment, they created datasets containing code with nonsensical or misleading variables, methods, and function names. Using CodeBERT for the analysis, they demonstrated that naming significantly affects LLMs ability to detect vulnerabilities, indicating that LLMs rely on clear and meaningful names.
Fang et al. [90] identified a gap in research, concerning how effective LLMs are at analyzing confusing, obfuscated, or poorly readable code. The authors found that GPT models were more likely than smaller models to generate accurate and detailed explanations for input code. On the other hand, obfuscated code impacted the ability of LLMs to generate explanations. GPT models had a drop in security analysis accuracy, while smaller models were unable to handle obfuscated code. In conclusion, GPT models excel in code analysis but struggle with obfuscated code, indicating that clarity of the code significantly impacts LLMs analysis performance.
@@ V-E LLMs Fine-tuning for Vulnerability Detection
To enhance the performance of LLMs on vulnerability detection tasks, several researchers have explored fine-tuning techniques. This process, illustrated in Figure 3, involves adapting pre-trained models to security-related tasks, by further training them on a security-specific dataset.
@@
After evaluating LLMs for vulnerability detection, the authors of [57] fine-tuned one of the models that performed poorly on vulnerability detection tasks, Meta-Llama-3-8B-Intruct, to verify the impact of fine-tuning on the domain-specific performance. The results showed that fine-tuning can have a positive impact on vulnerability detection tasks, but challenges in identifying exact vulnerabilities remain. Moreover, fine-tuning was also explored by Ding et al. [58]. Despite applying advanced techniques, such as class weights and contrastive learning, LLMs struggled with detecting realistic vulnerabilities.
@@ VI-A Uses of LLMs for Fixing Vulnerabilities
Berabi et al. [101] proposed a new method for fixing code vulnerabilities using LLMs, which enhances their ability to fix vulnerabilities. To increase the model performance, their approach combines static code analysis with code reduction, a technique that involves selecting only the code that contains a defect and the relevant context for interpreting it, and feeding both to the LLM. Results showed that this approach improved the efficiency of the model, by allowing it to understand and focus only on the sections that require fixing.
@@ VI-B LLMs Fine-tuning for Fixing Vulnerabilities
Wu et al. [105] evaluated the performance of different LLMs in fixing Java vulnerabilities, with and without fine-tuning, on two benchmarks they proposed. Across 5 non-fine-tuned LLMs, the best-performing model, Codex, fixed 20.4% of vulnerabilities. Concerning the fine-tuning experiment, the authors could fine-tune 4 models out of 5, since OpenAI does not provide a public API for fine-tuning Codex. Among the fine-tuned models, the best performer (i.e., fine-tuned InCoder), fixed only 18% of vulnerabilities, a small improvement of 8% over the non-fine-tuned version. Besides, these 4 fine-tuned models could only fix vulnerabilities that ureqired simple changes, such as deleting statements, renaming variables, or renaming methods. The models struggled with more complex vulnerabilities, such as CWE-325: Missing Cryptographic Step, and CWE-444: HTTP Request/Response Smuggling.
@@ VII LLMs for Detecting and Fixing Vulnerabilities in a Unified Process
In Section V and Section VI, we discussed using LLMs as tools for handling vulnerability detection and vulnerability fixing separately. In this section, we explore an integrated approach where LLMs are utilized to detect and fix security vulnerabilities in a unified process. Additionally, it examines the capability of LLMs to identify vulnerabilities in the code they generate, and automatically resolve them. Last, the section covers fine-tuning for enhancing LLMs performance, in a unified pipeline for vulnerability detection and fixing.